OpenStudy (anonymous):

(143(5) + 67)5 ___________ A. 1950 B. 1955 C. 1956 D. 1953

OpenStudy (tkhunny):

Please demonstrate your best efforts.

OpenStudy (tkhunny):

Is the last number supposed to be 2.5?

OpenStudy (anonymous):

no

OpenStudy (tkhunny):

Why would you respond to my second post and ignore the first? Show your work. You can't have nothing.

OpenStudy (anonymous):

y would i give the answer away when its for you guys to solve

OpenStudy (tkhunny):

If it is for us to solve, it is spam and you should receive a warning. If it is for you to learn, by first showing your best work, than that is the purpose of OpenStudy. You choose.

OpenStudy (anonymous):

Simplifying (143(5) + 67) * 5 = 0 Multiply 143 * 5 (715 + 67) * 5 = 0 Combine like terms: 715 + 67 = 782 (782) * 5 = 0 Multiply 782 * 5 3910 = 0 Solving 3910 = 0 Couldn't find a variable to solve for. This equation is invalid, the left and right sides are not equal, therefore there is no solution.

OpenStudy (just_one_last_goodbye):

Need help? ^_^

OpenStudy (anonymous):

yea but im pritty sure it B

OpenStudy (abhisar):

@aloud , there is no need of variable. You simply have to find the result.

OpenStudy (anonymous):

yea

OpenStudy (abhisar):

But I don't think any of the option match the correct answer.

OpenStudy (just_one_last_goodbye):

I got 3910 :/

OpenStudy (just_one_last_goodbye):

Let me try to simplify

OpenStudy (anonymous):

K

OpenStudy (anonymous):

im sorry i guess it is 3910

OpenStudy (tkhunny):

3910 is correct. Thus my question in the beginning, since 3910 / 2 = 1955 -- and that is in the answer list.